OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13

OpenStudy (anonymous):

x=-1, y=10/3

OpenStudy (anonymous):

we have many ways to start this problem ,the easiest one is adding Let x + 3y = 9 be equation 1 and let 3x – 3y = –13 be equation 2 By adding 1+2 u get : x+3y+3x-3y=9+-13 4x=-4 x=-1 now after u got the x value go and substitute in any equation and let it be equation 1 (x+3y=9) put x=-1 , -1+3y=9 , 3y=9+1 ,3y=10 ,y=10/3

OpenStudy (anonymous):

@phemo01 :Dont Give the final answer ,according to C.O.C u should help people to get their answer by their own by helping them and showing them steps for solving ..

OpenStudy (anonymous):

thank u guys